OpenStudy (lgbasallote):

let's see if you're as good as you claim to be @swissgirl Prove that there are no positive perfect cubes less 1000 that are the sum of the cubes of two positive integers

OpenStudy (lgbasallote):

Note: no brute force substitution allowed

OpenStudy (anonymous):

can i use fermat?

OpenStudy (lgbasallote):

....isn't that like 30 pages long?

OpenStudy (lgbasallote):

.well according to urban myth anyway....

OpenStudy (anonymous):

actually the proof for the cases \(n=4\) is not too bad, really like pythagoras the proof for the case \(n=3\) is also doable the general case is well... you know that story study for 20 years and then come back and explain it to me

OpenStudy (lgbasallote):

oh. so the one that's 30 pages long is a^n + b^n = c^n...well then how does n = 3 look like?